From GlueXWiki
Jump to: navigation, search

September 12, 2013 FDC meeting


  1. Installation preparations [1] (all)
  2. Engineering (Bill)
  3. Electronics (Chris, Nick)
  4. Tests in EEL126 (Beni, Vlad, Lubomir)
  5. Other


Participants: Bill, Dave, Vlad, Beni, Nick, and Lubomir.

Installation preparations

- As indicated by Elton, after the BCAL was inserted in the magnet, it might be down w.r.t. the magnet axis by ~1/8". BCAL will be surveyed and then we will have the final numbers. However they will start the installation of the rails inside BCAL soon and we discussed how they should be aligned. For the FDC the most important is to be aligned w.r.t. the beam line. We have copper-free area of 12mm radius and insensitive area of 30mm, so 3mm is significant given the sharp rise of the background rates with the distance to the beam. Beni: for the tracking it is important also to preserve the magnetic field symmetry, so the FDC should be aligned with the magnet axis. Hopefully, the magnet and the beam are well aligned.

- According to Bill the nominal spacer between the BCAL plate and the rails is 3/16" thick (4.76mm), which means one can still do adjustments of the rails if the BCAL offset is not more than 3mm. If it is more, it will mean FDC will sit asymmetrically w.r.t. the top and bottom rails and the complication might be the installation/access of the connectors against the rails.


- Bill made all the slots in the g10 skin pieces and cut the plastic holders. Slots have to be made in the holders too, and Tom will take care of this when he's back. Then we will install the skin on the FDC, however we have to remove it for the survey and then reinstall it.

- The long bars to be attached on the rails are ready and Bill coated them with Alodine. Few more parts are needed to install the bars.


- Chris fixed the pre-amp with bad pins (that was under the rail) and Vald installed and tested it. Vlad also swapped two pre-amps to have the correct labels. When removing/installing a pre-amp the bracket on the back side has to be kept with a tool. Before we used green epoxy to glue the brackets to the tube, but it was not so stable, and Dave and Bill think it is not worth doing it.

- Nick will give Beni at least 8 discriminator cards (pre-amps) for the tests with the new F1TDC.

Tests at EEL126

- Vlad finished the HV test of the FDC packages. The noise that we had in several channels disappeared after some HV training. Bill can go ahead with the installation of the skin.

- Beni got 8 new F1TDCs to be tested and started working on the DAQ software for them.

- The extra fADC125 (with double sample size) that was given to us by Cody and used to test the spare package showed a problem. It works OK at the beginning, but after some time it starts showing only big signals and the baseline is exactly at zero. Usually the first file is OK and at the beginning of the second file the data are corrupted. Once it happened also to all the fADCs but this might be another problem. Beni will discuss the problem with Cody. We will replace the module and check if the problem persists.